OpenStudy (anonymous):

Solve and express the solution set in simplest form. 4x-1/3 = 7/1 guys I really need help im so confused on this pls

OpenStudy (anonymous):

Well, the first thing we need to do is to put the x on a side of the equation all by itself. To do that we must add (1/3) to each side. At this point we have: 4x = 7/1 + 1/3

OpenStudy (anonymous):

Then, we need to add 7/1 and 1/3, because as they are, it's just too confusing. Well, we can multiply (7/1) by (3/3), to give us (21/3). So now, we have: 4x = 21/3 + 1/3

OpenStudy (anonymous):

And then all we've got to do is isolate the x by dividing our result by four.

OpenStudy (anonymous):

omg thank you so much ur my life saver ily